OpenStudy (anonymous):

Evaluate limit x->0 sin7x/sin4x

OpenStudy (anonymous):

7/4

OpenStudy (anonymous):

but how did you get that

OpenStudy (anonymous):

\[\text{Limit}\left[\frac{\text{Sin}[7x]}{\text{Sin}[4x]},x\to 0\right]=\frac{7}{4} \]